# The Value of Information in Human-AI Decision-making

### v4

@@ 3.2 Instance-level Agent-complementary Information Value
$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across all possible realizations defined by the data-generating model. To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal. This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions. Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal. To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$. Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v^{\prime}$ on instances with the signal realization $V=v$ is
$\mathrm{r}^{v,\pi,S}(v^{\prime})=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (7)
where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 3. If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
$\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (8)
Note that we allow $v^{\prime}\neq v$, i.e. the signal $v^{\prime}$ observed by the rational DM being different from the actual realization $v$. When $v^{\prime}\neq v$, the rational DM is misinformed, thus $\mathrm{r}^{v,\pi,S}(v^{\prime})\leq\mathrm{r}^{v,\pi,S}(v)$. $\operatorname{ILIV}$ in Definition 3.3 quantifies this improvement in payoff when the rational DM observes the actual $v$.
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v^{\prime}$ on instances where $V=v$ as:
$\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b})=\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b}).$ (9)
$\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the expected rational payoff on instances with $V=v$, where the rational DM only knows the information encoded in agent decisions $D^{b}$ without knowing any realizations of $V$. $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v,\pi,S}(v;D^{b})\geq\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
3.2.1 ILIV-SHAP Information-based Explanation
We apply $\operatorname{ILIV}$ to define an information-based explanation technique (ILIV-SHAP) that extends the canonical SHAP feature saliency explanation of a model’s prediction. ILIV-SHAP communicates how the data features lead to changes in the information value of AI predictions. While traditional SHAP summarizes the average contribution of each feature to a specific prediction over the baseline (prior) prediction, ILIV-SHAP summarizes the average contribution of each feature to the decision-relevant information value contained in the prediction.
Suppose a model $f$ that takes as input $m$ features and outputs a real number. Given an instance $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance of one feature $x_{i}$ to the model output $f(\mathbf{x})$ is encoded by the expected difference of model outputs when $x_{i}$ is marginalized out. Specifically, this is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$. Considering the interaction between features, SHAP (Lundberg and Lee, 2017) uses the Shapley value to quantify the importance scores averaged on different combinations of features:
$\phi_{i}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[g_{f}(\mathbf{x}^{\prime})-g_{f}(\mathbf{x}^{\prime}\backslash x_{i})]$
where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected model output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$. The scores $\phi_{i}(f,\mathbf{x})$ output by SHAP construct an explanation model for a model output, which quantifies the expected counterfactual change in the model output caused by the feature $x_{i}$.
ILIV-SHAP extends SHAP to give an explanation model of how features impact the decision-relevant information value of an individual model output.
@@ 3.2.1 ILIV-SHAP Information-based Explanation
$\phi_{i}^{\operatorname{ILIV}}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})]$
@@
where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized. This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum when $g_{f}(\mathbf{x}^{\prime})$ informs $f(\mathbf{x})$, i.e., $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$ – the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
The explanation model offered by ILIV-SHAP is grounded in the following two properties. First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$. Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017). Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$. Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. We demonstrate use of ILIV-SHAP in Appendix F.
3.3 Robust Analysis of Information Order
Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem. However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule. In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). Blackwell’s comparison of signals (Blackwell et al., 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks. (Li et al., 2022; Kleinberg et al., 2023).
Definition 3.5 (Blackwell Order of Information).
$\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
Definition 3.6.
Proposition 3.7 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$. Proposition 3.7 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
Proposition 3.7 (Hu and Wu 2024).
If $\forall\mu\in(0,1)$
$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
then $V_{1}$ is Blackwell more informative than $V_{2}$.
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix D.
4 Demonstrations
We demonstrate uses of our framework on three real-world decision-making tasks: medical diagnosis, deepfake video detection, and recidivism prediction.
First, we demonstrate the framework’s utility in model selection by evaluating how well different cardiac dysfunction prediction models trained on chest X-rays complement radiologists’ predictions (Rajpurkar et al., 2018; Johnson et al., 2019), showing how even among models with similar accuracy,some models strictly offer more complementary information than others over the radiologists. Next, we show how the framework can be used to evaluate how well humans exploit available features in a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022), how this compares to how an AI models exploits the information, and how providing AI assistance affects the human use. Suprisingly, even though AI assistance helps human-AI teams achieve higher decision accuracy, the $\operatorname{ACIV}$s of the video-level features did not change much compared to the $\operatorname{ACIV}$s over human-only predictions, indicating that simply displaying AI’s prediction did not help humans exploit the available information. Lastly, we demonstrate how ILIV-SHAP explanations surface unique insight into information value that humans lack relative to conventional SHAP for a recidivism prediction task (Angwin et al., 2022; Dressel and Farid, 2018). Detailed descriptions and results of all demonstrations are in Appendices E, D and F.
@@ 5 Experiment
    Does ILIV-SHAP improved human-AI decision-making over SHAP alone?
Our experiment assigns participants to one of two AI models with varying ex-ante ACIV and one of three explanation conditions: ILIV-SHAP and SHAP, SHAP, and no explanation. We hypothesize that the higher ACIV model will lead to better human-AI decisions and presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
Metric AI1 AI2
@@
Figure 1: Left: AI performance and ACIV. Right: Posterior predicted reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals.
@@
We study an AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al. (2022); Poursabzi-Sangdeh et al. (2021); Holstein et al. (2023)) . This task does not require any domain knowledge, making it suitable for a broad participant population. Additionally, the task is complex enough to benefit from AI assistance. We use the Ames, Iowa Housing Dataset (De Cock, 2011).
@@
We recruited 421 US-based participants via the crowdsourcing platform Prolific. Each participant was randomly assigned to 12 houses in the Ames Iowa Housing Dataset and performed a sequence of 24 trials (where two decisions were elicited for each house, one without AI assistance and one with), split evenly across a human-alone prediction phase and a human-AI prediction phase. In the first 12 trials, participants were asked to predict the house price of the assigned house using six features of the house, including year built, the living area above ground in square feet, the size of garage in car capacity, the number of fireplaces, year remodeled, and ratings of the house’s material and finish. We intentionally reduced the interpretability of the year remodeled and material and finish by witholding what variables they represented and rescaling their values. In the second 12 trials, participants were asked to revise their guesses in the first round after seeing the AI’s prediction and explanation. We incentivized participants with a base reward of $3.00$ USD and a bonus based on the mean squared error (MSE) between their guesses and the true house price in the 24 trials: $\textit{bonus}=3.00-\frac{\textit{MSE}}{3\times 10^{9}}$.
@@
Participants were assigned to one of six experimental conditions resulting from crossing the AI models (2 levels, where AI1 has high $\operatorname{ACIV}$ and AI2 has low $\operatorname{ACIV}$) with the explanation type (3 levels: ILIV-SHAP + SHAP, SHAP, and no explanation). We designed AI1 and AI2 so as to vary their potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features. On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add unique information over human judgments. We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1), so that model accuracy differences would not confound our results. In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction. To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same set-up prior to running the official study. We present ILIV-SHAP explanations of the six features that ranked and highlighted the features where the human-complementary information the AI’s prediction gets from the feature is higher than a threshold (see Appendix I for explanation examples).
@@
We report the $\operatorname{ACIV}$ of the AI models in the mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not. and the mean absolute percentage error (MAPE). To measure the human unassisted and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model [^4] https://aspredicted.org/jfm6-rrs3.pdf to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$ using R’s brms package (Bürkner, 2017).
[…]
@@
[…]

### v5

@@ 3.2 Instance-level Agent-complementary Information Value
$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across the distribution of all possible realizations defined by the data-generating model. To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal. This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from the access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions. Intuitively, this captures how much “room” there is for a specific signal to be better used. Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v^{\prime}$ on instances with the signal realization $V=v$ is
$\mathrm{r}^{v}(v^{\prime}):=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (6)
where $d^{r}(v^{\prime})$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 2. Note that we allow $v^{\prime}\neq v$, i.e. the signal $v^{\prime}$ observed by the rational DM can be different from the actual realization $v$. The expected payoff of a rational DM who is misinformed is guaranteed to be lower than is correctly informed, i.e., $\mathrm{r}^{v}(v^{\prime})\leq\mathrm{r}^{v}(v)$ for any $v^{\prime}$. This notion of flexibility allows to consider the information value of a counterfactual realization of a signal, which we use in designing the explanation of ILIV-SHAP in the next section.
If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
$\mathrm{r}^{v}(v^{\prime};D^{b}):=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (7)
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal realization $V=v^{\prime}$ on instances where $V=v$ as:
$\operatorname{ILIV}^{v}(v^{\prime};D^{b}):=\mathrm{r}^{v}(v^{\prime};D^{b})-\mathrm{r}^{v}(\emptyset;D^{b}).$ (8)
where $\mathrm{r}^{v}(\emptyset;D^{b})$ represents the expected rational payoff on instances of $V=v$, where the rational DM only knows the information encoded in agent decisions without knowing any realizations of $V$. $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v}(v;D^{b})\geq\operatorname{ILIV}^{v}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
Algorithm 2 A method to calculate the ILIV of $v^{\prime}$ on the instances of $v$.
1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, test counterfactual signal realization $v^{\prime}$, test signal realization $v$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
2:Output: $\operatorname{ILIV}^{v}(v^{\prime};D^{b})$
[…]
6:for $i=1$ to $k$ do
7:  $\hat{p}_{i}\leftarrow\hat{a}(v^{\prime},d^{b}_{j_{i}})$
8:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
9:  $\hat{p}^{b}_{i}\leftarrow\hat{a}^{b}(d^{b}_{j_{i}})$
10:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
11:  $s_{i}\leftarrow S(\hat{d}^{r}_{i},\omega_{i})-S(\hat{d}^{rb}_{i},\omega_{i})$
12:end for
13:return $\frac{1}{k}\sum_{i=1}^{k}s_{i}$
4 Information-based Explanation
We define an information-based explanation (ILIV-SHAP) to communicate where the AI prediction offers complementary information over the agent decisions. Traditional saliency-based explanations communicate the average contribution of each feature to a prediction over the baseline (prior) prediction, while ILIV-SHAP communicates the average contribution of each feature to the agent-complementary information value contained in the prediction.
Specifically, suppose a model $f$ that, for example, takes as input $m$ features and outputs a real number. Given an instance $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance of one feature $x_{i}$ to the model output $f(\mathbf{x})$ is encoded by the expected difference of model outputs when $x_{i}$ is marginalized out. This is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$. Considering the interaction between features, SHAP (Lundberg and Lee, 2017) uses the Shapley value to quantify the importance scores averaged on different combinations of features:
[…]
where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$ where $\mathbf{x}^{\prime}$ is the features that are not marginalized out.
@@ 4 Information-based Explanation
[…]
@@
where $\operatorname{ILIV}^{f(\mathbf{x})}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ over $D^{b}$ on the instances where the actual prediction is $f(\mathbf{x})$.
ILIV-SHAP shares the same properties as SHAP (Lundberg and Lee, 2017) by similarly constructing the importance scores. For example, ILIV-SHAP satisfies the critical efficiency axiom, i.e., the sum of the importance scores equals the information value of the model output, and the symmetry axiom, i.e., the importance scores are the same for any two features that contribute the same amount to the information value of the model output.
@@ 5 Experiment
    Does ILIV-SHAP improve human-AI decision-making over SHAP alone?
Our experiment assigns participants to one of two AI models with varying ex-ante ACIV {AI1=high ACIV, AI2=low ACIV}, and one of three explanation conditions {ILIV-SHAP and SHAP, SHAP, No Explanation}. We hypothesize that the higher ACIV model will lead to better human-AI decisions, and that presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
Metric AI1 (higher ACIV) AI2 (lower ACIV)
@@
Figure 1: Left: AI performance and ACIV. Right: Reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals according to posterior predictions of our regression model.
@@
We study AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al. (2022); Poursabzi-Sangdeh et al. (2021); Holstein et al. (2023)). This task does not require any domain knowledge, making it suitable for a broad participant population. Additionally, the task is complex enough to benefit from AI assistance. We use the Ames, Iowa Housing Dataset (De Cock, 2011).
@@
We recruited 421 US-based participants via the crowdsourcing platform Prolific. Each participant was randomly assigned to 12 houses out of 733 in the Ames Iowa Housing Dataset. Each completed a sequence of 24 trials, where two decisions were elicited for each house: one without AI assistance and one with. In the first 12 trials, participants were asked to predict the house price of the assigned house using six features of the house, including year built, the living area above ground in square feet, the size of garage in car capacity, the number of fireplaces, year remodeled, and ratings of the house’s material and finish. We intentionally reduced the interpretability of the year remodeled and material and finish by relabeling them as “Feature X” and “Feature Y” and rescaling their values. Therefore, the AI model that takes these two features as input is expected to have complementary information over human participants. In the second 12 trials, participants were asked to revise their guesses in the first round after seeing the AI’s prediction and explanation. We incentivized participants with a base reward of $3.00$ USD and a bonus based on the mean squared error (MSE) between their guesses and the true house price in the 24 trials: $\textit{bonus}=3.00-\frac{\textit{MSE}}{3\times 10^{9}}$.
@@
Participants were assigned to one of six experimental conditions resulting from crossing the two AI models with the three explanation types. We designed AI1 and AI2 to have varying potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features (which is meant to be bring down the accuracy of AI1 to be comparable with AI2). On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add complementary information over human judgments. We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1 left), so that model accuracy differences would not confound our results. In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction. To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same task and features prior to running the official study. We present ILIV-SHAP explanations of the six features, which ranked and highlighted the features where the human-complementary information (ILIV) over the AI’s prediction exceeds a threshold (see Appendix H for explanation examples).
@@
We report the $\operatorname{ACIV}$ of the AI models in mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not. and mean absolute percentage error (MAPE). To measure the human alone and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model with weakly informed priors444link to preregistration redacted for peer review to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$, using R’s brms package (Bürkner, 2017).
PE $\sim\text{student\_t}(\mu,\sigma,\nu)$
$\mu$ $=\textit{AI}*\textit{explanation}+\textit{round}+(1|\textit{participant\_id})$
$\log(\sigma)$ $=\textit{AI}*\textit{explanation}+\textit{round}+(1|\textit{participant\_id})$
@@
where AI and explanation are indicators of the experimental conditions, round is an indicator of whether the trial is in the first or second round, and participant_id is a unique identifier for each participant. We evaluate human-AI team performance by the reduction in absolute percentage error (APE) over human-alone predictions in the first round, i.e., $\text{Reduction in APE}=\text{\bf E}\![{\lvert\textit{PE}\rvert\mid\textit{round}=1}]-\text{\bf E}\![{\lvert\textit{PE}\rvert\mid\textit{round}=2}]$, where the expectation is taken over the Bayesian posterior distribution of the above Bayesian model.